OpenStudy (anonymous):

Help! How do you write equations of circles? A circle has a diameter with endpoints (3, -2) and (9, -8). What is the equation of the circle? r^2= (x+ 6)^2+ (y- 5)^2 r^2= (x+ 6)^2+ (y- 3)^2 r^2= (x- 6)^2+ (y+ 3)^2 r^2= (x- 6)^2+ (y+ 5)^2

OpenStudy (cwrw238):

first find the midpoint of the line joining the 2 points if the midpoint is (a,b) the equation is (x - a)^2 + (y - b)^2 = r^2

OpenStudy (cwrw238):

do you know the formula for the midpoint?

OpenStudy (anonymous):

Mid.= 3+9/2 , -2+-8/2 = 12/2 , -10/2 = 6,-5 = (x-6)^2 + (y+5)^2= R^2 Correct? Answer=D. ?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Thanks For Your Help! (:

OpenStudy (cwrw238):

yw